May you help me with this problem:
6x+2y=5
8x+2y=3
The instructions are to use linear combinations to solve the linear system.
Then check your solution. Thank you for time.

Answer by Alan3354(69443)

6x+2y=5
8x+2y=3
---------- Subtract
-2x = 2
x = -1
-----------
6x+2y=5
6(-1) + 2y = 5
2y = 11
y = 5.5